How many milliliters of 0.653 M NaOH are needed to neutralize 25.0 mL of a 1.02 M HBr solution? The neutralization reaction is: NaOH(aq) + HBr(aq) → H2O(l) + NaBr(aq)
1
Expert's answer
2019-05-21T05:54:36-0400

NaOH(aq) + HBr(aq) → H2O(l) + NaBr(aq)


25.0 mL of a 1.02M HBr solution contains "n(HBr)=c(HBr)\cdot V(HBr) = 1.02\frac{mmol}{mL}\cdot 25.0 mL = 25.5 mmol" HBr. According to stoichiometry of the reaction the same amount of NaOH is required for neutralisation of HBr:

"n(NaOH)=n(HBr)=25.5 mmol"

Volume of NaOH solution can be calculated:

"V(NaOH)=\frac{n(NaOH)}{c(NaOH)}=\frac{25.5mmol}{0.653\frac{mmol}{mL}}=39.1 mL"